A. Given the confidence interval for a population proportions is .0641<p<.0.863

express this in the form P^ + - E

B. Given that x= 60.7 and E=2.605

express the confidence interval for the population mean in the form X-E<u< X+E (ROUND TO ONE DECIMAL)
Answer #1

a)(note:I think interval is 0.641 < p< 0.863 that 0.0641 is very far, please revert)

here p=(0.863+0.641)/2=0.752

and E =(0.863-0.641)/2=0.111

hence P^ + - E   =0.752 -/+ 0.111

b)

confidence interval =60.7-2.605 < u <60.7+2.605 =58.095 < u <63.305